# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 3.2.4. Pluripotential theory and energy
using the non-pluripolar product introduced in [32], to define $\omega_{u_{\mu}}^{n}$. We will fix the additive constant by demanding that $\int u_{\mu}\omega_{0}^{n}=0$ and call $u_{\mu}$ the potential (function) of $\mu.$ Equivalently, the corresponding potential metric $\phi_{\mu}$ satisfies
@@
$\mathcal{J}(u)\coloneqq-\mathcal{E}(u)+\frac{1}{V}\int_{X}u\frac{\omega_{0}^{n}}{V},\,\,\,\,\,\mathcal{I}(u)\coloneqq\int_{X}u\frac{(\omega_{0}^{n}-\omega_{u}^{n})}{V}.$
@@
3.2.5. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
The space $\mathcal{H}(X,\omega_{0})$ of Kähler potentials can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to the space
@@ 3.2.5. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
in [40]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
@@ 3.2.7. The Futaki character of $(X,\Delta)$ and its quantization
More generally, 3.19 holds if $\psi$ and $\phi$ have finite energy and 3.20 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy. In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
@@
This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.19 first note that, for any given any $g\in\mathcal{G},$
@@
This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.5 defined by formula 3.5, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.19. Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$. Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e. iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.13, we can decompose
@@
Differentiating with respect to $t$ and using 3.4 thus proves 3.20. More generally, the proof of 3.19 only used that $\phi$ and $\psi$ have finite energy. Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [23] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [24], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy. Hence, $\frac{d}{dt}\mathcal{M}\left(\exp(t\text{Re}V^{1,0})^{*}\psi\right)_{|t=0}$, showing that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
The independence of the choice of $\phi_{0}$ is also a consequence of the fact that $\text{Fut}_{(X,\Delta)}$ arises as the leading term of the large-$k$ asymptotics of the quantized Futaki character $\text{Fut}_{(X,\Delta),k}(\cdot)$ at level $k,$ defined by
@@
where $\mathcal{L}_{V^{1,0}}$ now denotes the endomorphism of $H^{0}(X,-kK_{(X,\Delta)})$ induced by the infinitesimal action of $V^{1,0}$ on $H^{0}(X,-kK_{(X,\Delta)})$ (see the proof of [26, Prop 4.7]).

### v3

@@ 2.4. Gibbs polystability of log Fano curves
2.5. Higher dimensional examples
2.5.1. Gibbs stable orbifolds that are not exceptional
Following [34, 33], a log Fano variety $(X,\Delta)$ is said to be exceptional if $\text{lct}(X,\Delta;D)>1$ for any $\mathbb{Q}$-divisor $D$ linearly equivalent to $K_{(X,\Delta)}.$ By [33, Thm 1.7], this is equivalent to $\alpha(X,\Delta)>1,$ where $\alpha(X,\Delta)$ denotes Tian’s alpha invariant, whose algebro-geometric definition appears in formula 2.5. If $(X,\Delta)$ is exceptional, then $(X,\Delta)$ is Gibbs stable at any level $k.$ In fact, $(X,\Delta)$ is even uniformly Gibbs stable, by the lower bound 2.5.
Consider now the log pair $(X,\Delta)=(\mathbb{P}^{n},\Delta_{w}),$ where
$\Delta_{w}:=w(H_{1}+\cdots+H_{n+2}),\,\,\,w\in[0,1[$
for $(n+2)$ given hyperplanes $H_{i}$ in $\mathbb{P}^{n}$ with normal crossings. The log pair $(\mathbb{P}^{n},\Delta_{w})$ is Fano iff w<11/(n+2)w<1-1/(n+2) and exceptional iff w>11/n+1;w>1-1/n+1; a simple proof is provided in Remark 2.15 below.
Example 2.12.
When w=11/mw=1-1/m for $m\in\mathbb{Z}_{+}$ the corresponding orbifold $(\mathbb{P}^{n},\Delta_{w})$ is the global orbifold quotient of the Fermat hypersurface $X_{m}$ of degree $m$ in $\mathbb{P}^{n+1}$ under the standard action of the group $(\mathbb{Z}/\mathbb{Z}_{m})^{n}$ on $\mathbb{P}^{n+1},$ where $X_{m}$ is defined as the zero-locus of $x_{0}^{m}+\cdots+x_{n+1}^{m}$ in $\mathbb{P}^{n+1}.$ Indeed, $(\mathbb{Z}/\mathbb{Z}_{m})^{n}$ is the Galois group for the Galois branched cover $X_{m}\rightarrow\mathbb{P}^{n}$ defined by $(x_{0}:\cdots:x_{n+1})\mapsto(x_{0}^{m}:\cdots:x_{n}^{m}).$
The following result provides the first examples of Gibbs stable log Fano varieties – beyond complex log Fano curves – that are not exceptional:
Proposition 2.13.
The log Fano variety $(\mathbb{P}^{n},\Delta_{1-1/(n+1)}),$ which may be identified with the Fano orbifold $X_{n+1}/(\mathbb{Z}/\mathbb{Z}_{n+1})^{n},$ is Gibbs stable at any level $k.$
Although the group $\text{Aut}_{0}(\mathbb{P}^{n},\Delta)$ is trivial, the proof of the previous proposition exploits that $\text{Aut}_{0}(\mathbb{P}^{n})$ is non-trivial, by applying the following lemma to $X=\mathbb{P}^{n}$:
Lemma 2.14.
Let $(X,D)$ be a projective log pair and $G$ a connected algebraic group acting algebraically on $X.$ If $D$ is $G$-invariant, then so is the center $Z_{v}$ of any divisorial valuation $v$ over $X$ computing $\mathrm{lct}(X,D).$ In particular, given a Fano variety $X$ and a positive rational number $k$ such that $kK_{X}$ is defined as line bundle, any divisorial valuation computing the log canonical threshold of the $\mathbb{Q}$-divisor $\mathcal{D}^{(N_{k})}$ on $X_{N_{k}}$ is invariant under the diagonal action of $\text{Aut}_{0}(X).$
Proof.
Consider the action of $G$ on valuations $v$ over $X$ defined by
$(g\cdot v)(f)=v(g^{-1})f)$
for any $g\in G$ and rational function $f$ on $X.$ If $v$ is a divisorial valuation, then so is $g\cdot v.$ Indeed, if $v$ is represented by a prime divisor $E$ on a model $\pi:Y\rightarrow X,$ then $g\cdot v$ is represented by the same divisor $E$ wrt the model $g^{*}\pi:Y\rightarrow X.$ In particular, $A(v)=A(g\cdot v)$ and $Z_{gv}=g(Z_{v}).$ It follows that if $D$ is $G$-invariant and $v$ computes $\mathrm{lct}(X,D),$ then $g\cdot v$ also computes $\mathrm{lct}(X,D)$ and, as a consequence, $g(Z_{v})$ is the center of a valuation computing $\mathrm{lct}(X,D),$ for any $g\in G.$ But, for a general log pair $(X,D),$ the number of centers of valuations computing $\mathrm{lct}(X,D)$ is finite. Indeed, after scaling $D$ we may as well assume that $\mathrm{lct}(X,D)=1.$ Then, by definition, $v$ computes $\mathrm{lct}(X,D)$ iff $v$ is a log canonical place (in the terminology of the MMP) and it is well-known that there are only a finite number of centers in $X$ of log canonical places [1]. Finally, we can identify the center $Z_{v}$ with a point in the Chow variety $\mathcal{C}$ induced by a fixed ample line bundle on $X.$ The group $G$ naturally acts algebraically on $\mathcal{C}$ and since $G$ is connected so is the $G$-orbit $GZ_{v},$ viewed as a subset of $\mathcal{C}.$ But since the orbit $GZ_{v}$ is finite this implies that it consists of a single point, i.e. $Z_{v}$ is $G$-invariant, as desired. Finally, the last statement of the lemma follows from the fact that $\mathcal{D}^{(N_{k})}$ is is invariant under the diagonal action of $\text{Aut}_{0}(X)$ on $X^{N_{k}}$ (see the proof of Lemma 6.1). ∎
To prove Proposition 2.13 we start by expressing $K_{(\mathbb{P}^{n},\Delta_{w})}=:d\mathcal{O}(1),$ where d=n+1w(n+2)d=n+1-w(n+2) and $\Delta_{w}=w\Delta_{1}.$ The assumption w=11/(n+1)w=1-1/(n+1) is equivalent to
$d+w=1,$
since d+w=(n+1)(1w).d+w=(n+1)(1-w). Likewise, rescaling
$\mathcal{D}^{(N)}\eqqcolon dD_{N},\,\,\Delta_{N}=w\Delta_{1,N},$
where $\Delta_{1,N}$ is the divisor on $(\mathbb{P}^{n})^{N}$ defined by $\pi_{1}^{*}\Delta_{1}+\cdots+\pi_{N}^{*}\Delta_{1},$ in terms of the $N$ different projections $\pi_{i}\colon(\mathbb{P}^{n})^{N}\rightarrow\mathbb{P}^{n},$ we note that
(2.19) $\text{(i)\,}\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})=1,\,\,\,\text{(ii) }\mathrm{lct}((\mathbb{P}^{n})^{N},D_{N})\geq 1.$
The first item follows directly from the fact that $\Delta_{1,N}$ has simple normal crossings. The prove the second one observe that, by definition, $D_{N}$ is linearly equivalent to $\pi_{1}^{*}\mathcal{O}(1)+\cdots+\pi_{N}^{*}\mathcal{O}(1)$ and since $\mathrm{lct}((\mathbb{P}^{n},D)\geq 1$ for any $\mathbb{Q}$-divisor $D$ on $\mathbb{P}^{n}$ [34, Ex1.3] item $(ii)$ thus follows from formula 2.9. Now fix a divisorial valuation $v$ over $(\mathbb{P}^{n})^{N}$ and note that that the inverse of $\mathrm{lct}(X,D)_{v},$ appearing in the valuative formula 2.7 for $\mathrm{lct}(X,D),$ is linear wrt $D.$ Hence,
(2.20) $\frac{1}{\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)}+\Delta_{N}\right)_{v}}=d\frac{1}{\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right)_{v}}+w\frac{1}{\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})_{v}}\leq 1,$
giving $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)};\Delta\right)_{v}\geq 1.$ As a consequence, $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)};\Delta\right)\geq 1.$ Now consider a divisorial valuation $v$ over $(\mathbb{P}^{n})^{N}$ computing $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)}+\Delta_{N}\right).$ If $\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right)_{v}>1,$ then $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)}+\Delta_{N}\right)>1,$ by inequality 2.20. Consider finally the case when $\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right)_{v}=1.$ This implies, by item $ii$ in formula 2.19, that $v$ computes $\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right).$ To prove that $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)};\Delta\right)>1$ it will, by the inequality 2.20, be enough to show that $\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})_{v}>1.$ But, by the previous lemma, the center $Z_{v}$ of $v$ is invariant under group $\mathrm{PGL}(n+1,\mathbb{C})$ acting diagonally on $(\mathbb{P}^{n})^{N}.$ This, in fact, forces $v(\Delta_{1,N})=0$ and thus $\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})_{v}=\infty.$ Indeed, assume, to get a contradiction, that $v(\Delta_{1,N})>0.$ Then the center $Z_{v}$ is contained in support of $\Delta_{1,N}.$ In particular, fixing any $(x_{1},...,x_{N})\in Z_{v}$ this implies, since $Z_{v}$ is $\mathrm{PGL}(n+1,\mathbb{C})$-invariant, that $(gx_{1},...,gx_{N})$ is contained in $\Delta_{1,N}$ for any $g\in\mathrm{PGL}(n+1,\mathbb{C}).$ But this implies, since $\mathrm{PGL}(n+1,\mathbb{C})$ acts transitively on $\mathbb{P}^{n},$ that the support of the divisor $\Delta_{1}$ is all of $\mathbb{P}^{n},$ which is a contradiction.
Remark 2.15.
Taking $N=1$ and replacing the divisor $\mathcal{D}^{(N)}$ in the previous proof with any $\mathbb{Q}$-divisor that is linearly equivalent to $K_{(\mathbb{P}^{n},\Delta_{w})}$ shows that $(\mathbb{P}^{n},\Delta_{1-1(n+1)})$ is exceptional iff $d+w<1.$ Indeed, when $d+w<1,$ i.e. when w>11/n+1,w>1-1/n+1, the inequality 2.20 is strict and when $d+w\leq 1$ its left hand side is greater than are equal to $1.$ This is seen by taking $v$ to be the divisorial valuation on $\mathbb{P}^{n}$ defined by any of the hyperplanes $H_{i}.$
2.6. The minimal case on $\mathbb{P}^{n}$
As a preparation, we recall the following result [67, Theorem 11.2] describing the semistable locus $(\mathbb{P}^{n})_{\mathrm{ss}}^{N}$ of $(\mathbb{P}^{n})^{N}$ for general $N$ with respect to the diagonal action of $\mathcal{G}=\mathrm{PGL}(2,\mathbb{C})$ and polarization $K_{(\mathbb{P}^{n})^{N}}$: A point $(x_{1},\dots,x_{N})\in(\mathbb{P}^{n})^{N}$ is semistable iff for any proper linear subspace $W\in\mathbb{P}^{n},$
(2.21) $\{i:x_{i}\in W\}\leq\frac{\mathrm{dim}W+1}{n+1}N.$
With this explicit description we can compute $\gamma^{(N)}(\mathbb{P}^{n})$ for the minimal value of $N$.
Proposition 2.16.
$\mathbb{P}^{n}$ is Gibbs polystable at the minimal level $k_{\mathrm{min}}\coloneqq 1/(n+1)$, more precisely
$\gamma^{(N_{\mathrm{min}})}(\mathbb{P}^{n})^{\mathcal{G}}=\infty.$
Proof.
Consider $X=\mathbb{P}^{n}$ at the minimal level $k_{\min}\coloneqq 1/(n+1)$, i.e. $kK_{X}=\mathcal{O}(1)$ and $N_{\mathrm{min}}=n+1$. First recall the well-known fact that the quantized Futaki character of $\mathbb{P}^{n}$ vanishes at every level. Note that in homogeneous coordinates $X_{i,j}$, $1\leq i,j\leq N$ on $(\mathbb{P}^{n})^{n+1}$, $\mathrm{det}S^{(N_{\mathrm{min}})}=\mathrm{det}[X_{i,j}]_{i,j}$. Note that when $N=n+1$, condition 2.21 for $W$ of codimension 1 states that not all $x_{i}$ lie on $W$, or equivalently that $\mathrm{det}S^{(n+1)}(x_{1},\dots,x_{n+1})\neq 0$. Thus $(\mathbb{P}^{n})_{\mathrm{ss}}^{N_{\mathrm{min}}}\subset\{\mathbb{\mathrm{det}}S^{(N_{\mathrm{min}})}\neq 0\}$ so we get
$\gamma^{(N_{\mathrm{min}})}(\mathbb{P}^{n})^{\mathcal{G}}=\mathrm{lct}((\mathbb{P}^{n})_{\mathrm{ss}}^{N_{\mathrm{min}}},\frac{1}{k_{\mathrm{min}}}\mathrm{det}S^{N_{\min}})=\infty.$
In fact, when $N=n+1$ in 2.21, the conditions for arbitrary $W$ follows from the ones for $W$ of codimension 1 and thus $(\mathbb{P}^{n})_{\mathrm{ss}}^{N_{\mathrm{min}}}=\{\mathbb{\mathrm{det}}S^{(N_{\mathrm{min}})}\neq 0\}$ in this special case.
By Theorem 3.13, $\gamma^{(N)}(X)^{\mathcal{G}}$ is in not increasing, in general. This phenomenon persists for $X=\mathbb{P}^{n}$ when $n\geq 2$ by the following observation. Consider $X=\mathbb{P}^{2}$ at level $k=\frac{2}{3}$ so that $kK_{\mathbb{P}^{2}}=\mathcal{O}(2)$. Then it does not hold that $(\mathbb{P}^{2})_{\mathrm{ss}}^{6}\subset\{\mathrm{det}S^{(6)}\neq 0\}$. To see this, pick $6$ distinct points $(y_{1},\dots,y_{6})$ on a smooth quadric in $\mathbb{P}^{2}$, then $\mathrm{det}S^{(6)}(y_{1},\dots,y_{6})=0$. Since they are distinct, the only condition to be checked for semistability is whether $5$ or more points are contained in a single line. But a line meets a quadric in $\mathbb{P}^{2}$ in at most $2$ points by Bezout, so at most two of the $y_{1},\dots,y_{6}$ are contained in a common line. By 2.21, $(y_{1},\dots,y_{6})$ is semistable, giving
$\mathrm{lct}_{(y_{1},\dots,y_{6})}((\mathbb{P}^{2})^{6},\frac{3}{2}\mathrm{det}S^{(6)})\leq 2/3$
and thus $\gamma^{(6)}(\mathbb{P}^{2})\leq 2/3$ and $\mathbb{P}^{2}$ is not Gibbs polystable at level $k=2/3$.
@@ 3.2.4. Pluripotential theory and energy
using the non-pluripolar product introduced in [38] to define $\omega_{u_{\mu}}^{n}$. We will fix the additive constant by demanding that $\int u_{\mu}\omega_{0}^{n}=0$ and call $u_{\mu}$ the potential (function) of $\mu.$ Equivalently, the corresponding potential metric $\phi_{\mu}$ satisfies
@@
$\mathcal{J}(u)\coloneqq-\mathcal{E}(u)+\int_{X}u\frac{\omega_{0}^{n}}{V},\,\,\,\,\,\mathcal{I}(u)\coloneqq\int_{X}u\frac{(\omega_{0}^{n}-\omega_{u}^{n})}{V}.$
@@
3.2.5. The actions of $\mathcal{G}$ on $\text{$\text{PSH}$(}L),$ $\text{PSH}(X,\omega_{0})$ and $\mathcal{P}(X)$
The group $\mathcal{G}$ naturally acts from the right on $\text{$\text{PSH}$(}L)$ by pullback: $\phi\cdot g:=g^{*}\phi.$ Fixing a reference metric $\phi_{0}$ in $L$ and setting $\omega_{0}:=dd^{c}\phi_{0},$ the right action of $\mathcal{G}$ on $\text{$\text{PSH}$(}L)$ induces a right action on $\text{PSH(}(X,\omega_{0}),$ by identifying $\text{$\text{PSH}$(}L)$ with $\text{PSH}(X,\omega_{0}).$ In other words, $u\cdot g:=g^{*}(\phi_{0}+u)-\phi_{0}$ (which is thus different than the action of $\mathcal{G}$ on $\text{PSH(}(X,\omega_{0})$ by pullback). The group $\mathcal{G}$ also naturally acts on the right on $\mathcal{P}(X)$ by setting $\mu\cdot g:=(g_{-1})_{*}\mu:=g^{*}\mu.$ This action is compatible with the action of $\mathcal{G}$ on $\text{$\text{PSH}$(}L)$ in the following sense: if $\mu=(dd^{c}\phi)^{n}/V,$ then $g^{*}\mu=(dd^{c}g^{*}\phi)^{n}/V.$ Moreover, if $L=-K_{(X,\Delta)}$ then $g^{*}\mu_{\phi}=\mu_{g^{*}\phi}.$ Finally, there is also a natural left action of $\mathcal{G}$ on $\text{$\text{PSH}$(}L),$ defined by $g\cdot\phi:=(g^{-1})^{*}\phi,$ which is compatible with the natural left action of $\mathcal{G}$ on $\mathcal{P}(X),$ defined by $g\cdot\mu:=g_{*}\mu.$
3.2.6. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$ and the $L^{1}$-Finsler completion of $\mathcal{H}(X,\omega)$
The space $\mathcal{H}(X,\omega_{0})$ of Kähler potentials can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the length of a tangent vector $\dot{u}$ at $u$ is defined by
(3.13) $\left\|\dot{u}\right\|_{u}:=\left(\int_{X}\dot{u}^{2}\omega_{u}^{n}\right)^{1/2}$
The corresponding notion of geodesics was extended to the space
@@
in [51]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$ More generally, following [59], fixing $p\in[1,\infty[$ and replacing the $L^{2}$-norm in formula 3.13 with the corresponding $L^{p}$-norm yiels a Finsler metric on $\mathcal{H}(X,\omega_{0}),$ which induces a metric $d_{p}$ on $\mathcal{H}(X,\omega_{0}).$ As shown in [59], for $p=1,$ the metric space completion $\overline{(\mathcal{H}(X,\omega_{0}),d_{1})}$ of $(\mathcal{H}(X,\omega_{0}),d_{1})$ may be identified with $\mathcal{E}^{1}(X,\omega)$ (defined in the previous section). Thus $\overline{(\mathcal{H}(X,\omega_{0}),d_{1})}$ contains $\mathcal{H}_{1,1}(X,\omega)$ and, as shown in [59], the geodesics in $\mathcal{H}_{1,1}(X,\omega),$ discussed above, yield constant speed geodesics in the metric space $\overline{(\mathcal{H}(X,\omega_{0}),d_{1})}$. Moreover, on the subset $\{\mathcal{E}(u)=0\}$ of $\mathcal{E}^{1}(X,\omega)$ the following inequalities hold:
for some positive constants $A$ and $B.$ Indeed, by 3.11, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [11, Lemma 2.13] and by [61, Prop 5.5] the inequalities 3.14 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u)$ (see also [61, Remark 3.9]).
@@ 3.2.8. The Futaki character of $(X,\Delta)$ and its quantization
More generally, 3.21 holds if $\psi$ and $\phi$ have finite energy and 3.22 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy. In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
@@
This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.21 first note that, for any given any $g\in\mathcal{G},$
@@
This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.5 defined by formula 3.5, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.21. Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$. Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e. iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.15, we can decompose
@@
Differentiating with respect to $t$ and using 3.4 thus proves 3.22. More generally, the proof of 3.21 only used that $\phi$ and $\psi$ have finite energy. Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [25] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [26], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy. Hence, $\frac{d}{dt}\mathcal{M}\left(\exp(t\text{Re}V^{1,0})^{*}\psi\right)_{|t=0}$, showing that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
The independence of the choice of $\phi_{0}$ is also a consequence of the fact that $\text{Fut}_{(X,\Delta)}$ arises as the leading term of the large-$k$ asymptotics of the quantized Futaki character $\text{Fut}_{(X,\Delta),k}(\cdot)$ at level $k,$ defined as follows when $k$ is a positive integer such that $k\Delta$ is divisor:
@@
where $\mathcal{L}_{V^{1,0}}$ now denotes the endomorphism of $H^{0}(X,-kK_{(X,\Delta)})$ induced by the infinitesimal action of $V^{1,0}$ on $H^{0}(X,-kK_{(X,\Delta)})$ (see the proof of [28, Prop 4.7]). More generally, when it is only assumed that $kK_{(X,\Delta)}$ is a line bundle we define $\text{Fut}_{(X,\Delta),k}$ by formula 6.1 in Lemma 6.1.